OpenStudy (anonymous):

Use the rule y = -x + 7 to fill in the blank. If (-1, y) is on the graph, then y =

OpenStudy (alexandervonhumboldt2):

this is easy: substitute -1 for x and solve for y. y=-(-1)+7=1+7=???

OpenStudy (anonymous):

8?

OpenStudy (alexandervonhumboldt2):

yeah